OpenStudy (howard-wolowitz):

Can someone help me here:

OpenStudy (howard-wolowitz):

can someone show me the set-up version of this problem

OpenStudy (yanasidlinskiy):

Are you asking like for an example? A = P(1 + r/n)^nt where: A = the accrued amount P = the original number r = rate of increase as a decimal n = number of times compounded per year (will be 1 in this case) t = time in years

OpenStudy (anonymous):

Assume that both of problems follow exponential growth/decay rule with general function is: y=y(0)ekt with k is growth/decay constant rate. (1) Deer population function y(0) = 1253 k = 3% t = 6 (years) So y(6) = 1253 . e^(0.03*6) ~ 1500 (deers) The answer is approximated, so the most close correct option is C: 1496 deers (2) Growth factor b is rate 9.5% = 0.095 (A)